OpenStudy (anonymous):

MEDAL REWARDED! Just two questions please help! Evaluate the following put in simplest form no decimals 1. (9/16)^3/2 2. (0.6^-2)^-1/2

OpenStudy (mathmale):

Regarding (9/16)^3/2 : This can be re-written, in the Equation Editor, as \[(\frac{ 9 }{ 16 })^{3/2}\]

OpenStudy (mathmale):

Note that the exponent 1/2 denotes square root, and that 3/2 can be re-written as (1/2)^3. Then \[(\frac{ 9 }{ 16 })^{3/2}=[(\frac{ 9 }{ 16 })^{1/2}]^{3}\]

OpenStudy (mathmale):

What would you do next, to further simplify this?

OpenStudy (anonymous):

exponent form? I'm not quite sure to be honest